Substitution method
y=-3x+13
2x+9y=-8

Answer by Fombitz(32388)

2x+9y=-8
2x+9(-3x+13)=-8
2x-27x+117=-8
-25x=-125
x=5
y=-3x+13
y=-3(5)+13
y=-2
(5,-2)